# How to Securely Shuffle? A survey about Secure Shufflers for privacy-preserving computations

### v1

@@ 5.4. Towards new shuffling-based applications
For example, Wang et al. (2025) recently proposed a shuffling-based taxi-hailing service. In this application, users willing a taxi can submit their request and location through a secure shuffler and the secure shuffler will provide DP guarantees on their location data. More generally, the authors introduced the notion of “private individual computations:” a broad range of “permutation-equivariant” computations such as taxi-hailing and information retrieval. In such protocols, each user receives a personalized service/output, while secure shufflers amplify the privacy. Thus, their work opened the way to new applications of secure shufflers.
Another recent example is the shuffling-based private information retrieval described by (Gascón et al., 2024).
@@ 5.5. Applications unrelated to privacy-preserving computations
Our survey focuses specifically on application to privacy-preserving computations as they are the protocols that revived the interest in secure shufflers. However, there exist other “historical” applications that we do not survey: web anonymization (Dingledine et al., 2004), anonymous messaging (Sasy and Goldberg, 2023), electronic voting (Neff, 2001; Juels et al., 2010), and blockchain operations (Ruffing et al., 2014).
@@ 6.1.7. Questions 7: Are the data owners and/or the servers trusted for availability?
Possible answers: Yes or No (×$\times 2$).
@@ 6.2.1. Medical surveys
A group of hospitals wants to deploy a privacy-preserving surveying system. Each hospital agreed to deploy a shuffling server. The data owners (i.e., survey participants) answer the surveys on their mobile phones and submit their answers through the shuffler. We suppose the practitioner wants to use shuffling-based data (Ishai et al., 2006) aggregation to preserve privacy.
@@ 6.2.3. Smart keyboard application
This last use case was used as motivation in the paper that sparked the interest in FL (Konečný et al., 2016): a Google-like company develops a smartphone keyboard application and wants to improve the next-word recommendations. Each smartphone locally trains a model based on the application usage, and a central server securely aggregates the local models to produce a global recommendation model. We want a single-server secure shuffler to perform the secure aggregation step because the company does not want to involved any third party. Google-like services have millions (or even billions) of users, so the scalability expectations are high.
@@ 7. Perspectives on the literature
This section introduces several general perspectives on the literature. The subsection are independent from each other and discuss misconceptions or subtle details that are often overlooked in the literature. Moreover, these perspectives enable to identify several promising future works.
@@ 7.1. Preventing metadata leakage is unnecessary
Our work focused on the most conventional definition of sender anonymity, but some works also proposed stronger anonymity notions than Definition 3.4: referred to as ultimate anonymity (Gelernter and Herzberg, 2013) or perfect anonymity (Backes et al., 2013). This stronger anonymity forbids the metadata leakage mentioned in Section 3: the size and number of sent messages.
From a theoretical point of view, it combines sender anonymity with a notion called “unobservability”. Unobservability ensures that the observation of the traffic leaks absolutely no information. Gelernter and Herzberg (Gelernter and Herzberg, 2013) show that ultimate anonymity is inefficient because each sender must generate a large amount of cover traffic to prevent metadata leakage. Indeed, each sender must send $O(n)$ data to be a possible sender of the whole batch, even if the sender has small or no data to send. This result explains why Dissent (the only secure shuffler satisfying ultimate anonymity) is the least scalable technique.
@@
On the one hand, privacy-preserving computations usually involve messages containing fixed-sized vectors or atomic elements (e.g., floating-point numbers). Thus, they are all the same size. On the other hand, many privacy-preserving computations define the number of messages sent by each data owner independently of their private data. For example, in shuffling-based secure aggregation (Ishai et al., 2006), each data owner must send the same number of messages. To sum up, (shuffling-based) privacy-preserving computations do not require anonymity stronger than the sender anonymity defined in Section 3 because the message metadata is often fixed and known publicly.
@@ 7.2. Imperfect shuffling/anonymity can be acceptable
Shuffling-based private computations (including in DP) usually assume a perfect shuffle (i.e., uniform anonymity distribution as defined in Section 3.3.1). However, Ishai et al. (2006) claimed that only a “sufficient uncertainty” was necessary; i.e., imperfect shuffle could be acceptable. Unfortunately, they provided no concrete results supporting this claim. Recently, Balle et al. (2023) showed that this intuition was true as their results holding also for an imperfect shuffle.
@@ 7.3. Avoid modifying the shuffler functionality
Even though shuffle DP works agree on the intuitive definition of a shuffler, there is no consensus regarding the exact capability of a shuffler. Most papers (Cheu et al., 2019; Cheu, 2021) rely on the most basic definition as we do (i.e., the shuffler simply outputs a permutation), but some papers (Yang et al., 2023) complete this functionality with other simple tasks such as adding unique identifiers to the messages.
@@ 7.4. Information-theoretic Shuffle DP is impractical
Computational DP (Beimel et al., 2008; Mironov et al., 2009) is a relaxed variant of DP where the DP guarantees hold against computationally-bounded adversaries. This definition is opposed to “information-theoretic” DP (i.e., the initial DP definition proposed by Dwork (Dwork, 2006)), where the guarantees also hold against computationally unbounded adversaries. This separation is classic in security, where we oppose computational security to information-theoretic security.
A DP algorithm is “computationally differentially private” if one of its components is only computationally secure. For example, as mentioned in (Bun et al., 2016), information-theoretic DP does not permit using cryptographically secure pseudo-random generators in place of perfect randomness.
This constraint (i.e., no computationally secure component) is problematic for Shuffle DP. As mentioned in Section 4, all secure shufflers assume the existence of a secure communication channel between the senders and the shuffler. However, Maurer (Maurer, 1999) showed that information-theoretically secure communication channels are impractical. Therefore, we should only assume computationally secure communication channels in the context of shuffle DP. Hence, only computational shuffle DP can be implemented efficiently.
Relying on “information-theoretic” shuffle DP is then impractical because any real world deployment requires a computational hardness assumption. Thus, shuffle DP applications should take this mandatory assumption into consideration in their analysis. Such consideration should incentivize more “computational” DP results; known to require less noise than information-theoretic DP (Mironov et al., 2009).
@@ 7.5. Malicious data owners are a serious threat
Our security properties hold against an unbounded number of malicious data owners. Anonymous communications systems commonly assume a “sufficiently high” number of honest senders, as anonymity holds within the set of honest senders. For example, Langowski et al. (Langowski et al., 2023) highlighted this limitation when presenting their anonymous broadcaster.
@@
To verify that malicious data owners add an appropriate noise, Biswas and Cormode (Biswas and Cormode, 2022) introduced the notion of verifiable DP (generalized to Shuffle DP by Bontekoe et al. (2024)); leveraging zero-knowledge proofs to verify the DP noise generation and addition.
@@
Malicious data owners also threaten the computation correctness because they can bias the final results via data poisoning attacks. For example, in FL, this takes the form of model poisoning attacks (Cao et al., 2021; Carlini, 2021; Schuster et al., 2021) in which attackers submitting local models that would introduce undesired behaviors in the global model. Such poisoning can have serious consequences as it can, for example, add a backdoor to an ML model (Schuster et al., 2021).
Norm bounding is a popular mitigation technique (Shejwalkar et al., 2021). This simple mitigation has been incorporated into secure aggregation protocols using ZKPs (Bell et al., 2022b; Chowdhury et al., 2022).
However, no mitigation compatible with shuffling-based protocols has been yet described. This verification is trivial when each data owner sends a single message because the verification can be done on the shuffle output. However, in multi-message Shuffle DP (Cheu, 2021) or shuffling-based secure aggregation (Ishai et al., 2006), each data owner splits its private value into several messages. A zero-knowledge verification is necessary to verify claims about the initial private value while protecting anonymity.
@@ 7.6. Trusted hardware is not a silver bullet
In theory, trusted hardware (i.e., TEE) is an appealing assumption that can solve privacy issues: the protocol is secure as long as the data goes through the trusted hardware (even with a malicious server owner). Several papers promoted TEE in FL (Bittau et al., 2017; Boutet et al., 2021; Zhang et al., 2021; Zhao et al., 2022) motivated by its performances and (theoretical) security. However, one must also understand the strength of the assumption required by TEEs: the data owner fully trusts the manufacturer. TEE also requires trusting the protocol implementation because there exist other flaws: side-channel attacks (Brasser et al., 2017; Wang et al., 2017).
Sasy et al. (Sasy et al., 2022) presented an excellent example of the risk of TEE. While some shuffle DP papers used Prochlo (Bittau et al., 2017) (a TEE-based secure shuffler) as motivation, Sasy et al. (Sasy et al., 2022) argued that recent attacks against TEE make Prochlo insecure. Hence, even if Prochlo seemed at first totally secure, it turned out to be insecure in practice due to TEE vulnerabilities.
@@ 8. Conclusion and future directions
Linear complexity for 1-server secure shuffling. Recent works (Bittau et al., 2017; Shi and Wu, 2021; Bünz et al., 2021) approached linear complexity (i.e., $O(n)$ communications and computations) for 1-server secure shuffling, the optimal complexity for 1-server shuffling.
Prochlo has linear complexity but is vulnerable to side-channel attacks. Recent secure TEE-based solutions (Ngai et al., 2024; Gu et al., 2023) avoided these attacks, but have a complexity of $O(n\log n)$. Finally, NIDAR verifies fewer security properties than the rest of the solutions and only has subquadratic costs.
@@
Formalization of the “imperfect” shuffle DP model. Several works (Kwon et al., 2017; Bünz et al., 2021; Langowski et al., 2023; Zhou and Shi, 2022; Balle et al., 2023) proposed relaxed anonymity definitions to conceive more scalable secure shufflers. While some works (Zhou and Shi, 2022; Balle et al., 2023) adapted the shuffle DP model to their “imperfect” shuffling, no DP results exist for the relaxation of Langowski et al. (Langowski et al., 2023). Thus, further theoretical work is still needed to formalize these relaxations, and present DP results for shufflers with a relaxed anonymity definition.
Computational shuffle DP. Subsection 7.4 explained why only computational Shuffle DP is practical. While Mironov et al. (Mironov et al., 2009) showed that computational DP can reach better accuracy guarantees than information-theoretic DP, no work has yet studied whether this relaxation in Shuffle DP. Hence, analogous investigations for Shuffle DP would be necessary to take advantage of our impossibility result and reach higher accuracy guarantees.
New types of shuffling-based private computation. While most existing applications are using secure shufflers to perform a form of data aggregation, recent works (Gascón et al., 2024; Wang et al., 2025) recently opened the way to other privacy-sensitive applications such as information retrieval or taxi hailing could be enabled by secure shufflers. Thus, further research into new shuffling-based private computations is a promising direction. The concept of “private individual computation” introduced by Wang et al. (2025) represents a valuable framework to structure such research effort.

### v2

@@ 5.4. Towards new shuffling-based applications
For example, wang_beyond_2025 recently proposed a shuffling-based taxi-hailing service. In this application, users willing a taxi can submit their request and location through a secure shuffler and the secure shuffler will provide DP guarantees on their location data. More generally, the authors introduced the notion of “private individual computations:” a broad range of “permutation-equivariant” computations such as taxi-hailing and information retrieval. In such protocols, each user receives a personalized service/output, while secure shufflers amplify the privacy. Thus, their work opened the way to new applications of secure shufflers.
Another recent example is the shuffling-based private information retrieval described by (gascon_computationally_2024).
@@ 5.5. Applications unrelated to privacy-preserving computations
Our survey focuses specifically on application to privacy-preserving computations as they are the protocols that revived the interest in secure shufflers. However, there exist other “historical” applications that we do not survey: web anonymization (dingledine2004tor), anonymous messaging (sasy_sok_2023), electronic voting (neff_verifiable_2001; juels_coercion-resistant_2010), and blockchain operations (kutylowski_coinshuffle_2014).
@@ 6.1.7. Questions 7: Are the data owners and/or the servers trusted for availability?
Possible answers: Yes or No ($\times 2$).
@@ 6.2.1. Medical surveys
A group of hospitals wants to deploy a privacy-preserving surveying system. Each hospital agreed to deploy a shuffling server. The data owners (i.e., survey participants) answer the surveys on their mobile phones and submit their answers through the shuffler. We suppose the practitioner wants to use shuffling-based data (ishai_cryptography_2006) aggregation to preserve privacy.
@@ 6.2.3. Smart keyboard application
This last use case was used as motivation in the paper that sparked the interest in FL (konecny_federated_2016): a Google-like company develops a smartphone keyboard application and wants to improve the next-word recommendations. Each smartphone locally trains a model based on the application usage, and a central server securely aggregates the local models to produce a global recommendation model. We want a single-server secure shuffler to perform the secure aggregation step because the company does not want to involved any third party. Google-like services have millions (or even billions) of users, so the scalability expectations are high.
@@ 7. Perspectives on the literature
This section introduces several general perspectives on the literature. The subsections are independent of each other and discuss misconceptions or subtle details that are often overlooked in the literature. Moreover, these perspectives enable to identify several promising future works.
@@ 7.1. Preventing metadata leakage is unnecessary
Our work focused on the most conventional definition of sender anonymity, but some works also proposed stronger anonymity notions than Definition 3.4: referred to as ultimate anonymity (gelernter_limits_2013) or perfect anonymity (backes_anoa_2013). This stronger anonymity forbids the metadata leakage mentioned in Section 3: the size and number of sent messages.
From a theoretical point of view, it combines sender anonymity with a notion called “unobservability”. Unobservability ensures that the observation of the traffic leaks absolutely no information. Gelernter and Herzberg (gelernter_limits_2013) show that ultimate anonymity is inefficient because each sender must generate a large amount of cover traffic to prevent metadata leakage. Indeed, each sender must send $O(n)$ data to be a possible sender of the whole batch, even if the sender has small or no data to send. This result explains why Dissent (the only secure shuffler satisfying ultimate anonymity) is the least scalable technique.
@@
On the one hand, privacy-preserving computations usually involve messages containing fixed-sized vectors or atomic elements (e.g., floating-point numbers). Thus, they are all the same size. On the other hand, many privacy-preserving computations define the number of messages sent by each data owner independently of their private data. For example, in shuffling-based secure aggregation (ishai_cryptography_2006), each data owner must send the same number of messages. To sum up, (shuffling-based) privacy-preserving computations do not require anonymity stronger than the sender anonymity defined in Section 3 because the message metadata is often fixed and known publicly.
@@ 7.2. Imperfect shuffling/anonymity can be acceptable
Shuffling-based private computations (including in DP) usually assume a perfect shuffle (i.e., uniform anonymity distribution as defined in Section 3.3.1). However, ishai_cryptography_2006 claimed that only a “sufficient uncertainty” was necessary; i.e., imperfect shuffle could be acceptable. Recently, balle_amplification_2023 showed that this intuition was true as their results holding also for an imperfect shuffle.
@@ 7.3. Avoid modifying the shuffler functionality
Even though shuffle DP works agree on the intuitive definition of a shuffler, there is no consensus regarding the exact capability of a shuffler. Most papers (cheu_distributed_2019; cheu_differential_2021) rely on the most basic definition as we do (i.e., the shuffler simply outputs a permutation), but some papers (yang_adastopk_2023) complete this functionality with other simple tasks such as adding unique identifiers to the messages.
@@ 7.4. Information-theoretic Shuffle DP is impractical
Computational DP (beimel_distributed_2008; mironov_computational_2009) is a relaxed variant of DP where the DP guarantees hold against computationally-bounded adversaries. This definition is opposed to “information-theoretic” DP (i.e., the initial DP definition proposed by Dwork (dwork_differential_2006)), where the guarantees also hold against computationally unbounded adversaries. This separation is classic in security, where we oppose computational security to information-theoretic security.
A DP algorithm is “computationally differentially private” if one of its components is only computationally secure. For example, as mentioned in (bun_separating_2016), information-theoretic DP does not permit using cryptographically secure pseudo-random generators in place of perfect randomness.
This constraint (i.e., no computationally secure component) is problematic for Shuffle DP. As mentioned in Section 4, all secure shufflers assume the existence of a secure communication channel between the senders and the shuffler. However, Maurer (maurer_information-theoretic_1999) showed that information-theoretically secure communication channels are impractical. Therefore, we should only assume computationally secure communication channels in the context of shuffle DP. Hence, only computational shuffle DP can be implemented efficiently.
Relying on “information-theoretic” shuffle DP is then impractical because any real world deployment requires a computational hardness assumption. Thus, shuffle DP applications should take this mandatory assumption into consideration in their analysis. Such consideration should incentivize more “computational” DP results; known to require less noise than information-theoretic DP (mironov_computational_2009).
@@ 7.5. Malicious data owners are a serious threat
Our security properties hold against an unbounded number of malicious data owners. Anonymous communications systems commonly assume a “sufficiently high” number of honest senders, as anonymity holds within the set of honest senders. For example, Langowski et al. (langowski_trellis_2023) highlighted this limitation when presenting their anonymous broadcaster.
@@
To verify that malicious data owners add an appropriate noise, Biswas and Cormode (biswas_verifiable_2022) introduced the notion of verifiable DP (generalized to Shuffle DP by bontekoe_efficient_2024); leveraging zero-knowledge proofs to verify the DP noise generation and addition.
@@
Malicious data owners also threaten the computation correctness because they can bias the final results via data poisoning attacks. For example, in FL, this takes the form of model poisoning attacks (cao_data_2021; carlini_poisoning_2021; schuster_you_2021) in which attackers submitting local models that would introduce undesired behaviors in the global model. Such poisoning can have serious consequences as it can, for example, add a backdoor to an ML model (schuster_you_2021).
Norm bounding is a popular mitigation technique (shejwalkar_back_2021). This simple mitigation has been incorporated into secure aggregation protocols using ZKPs (bell_acorn_2022; chowdhury_eiffel_2022).
However, no mitigation compatible with shuffling-based protocols has been yet described. This verification is trivial when each data owner sends a single message because the verification can be done on the shuffle output. However, in multi-message Shuffle DP (cheu_differential_2021) or shuffling-based secure aggregation (ishai_cryptography_2006), each data owner splits its private value into several messages. A zero-knowledge verification is necessary to verify claims about the initial private value while protecting anonymity.
@@ 7.6. Trusted hardware is not a silver bullet
In theory, trusted hardware (i.e., TEE) is an appealing assumption that can solve privacy issues: the protocol is secure as long as the data goes through the trusted hardware (even with a malicious server owner). Several papers promoted TEE in FL (bittau_prochlo_2017; boutet_mixnn_2021; zhang_shufflefl_2021; zhao_sear_2022) motivated by its performances and (theoretical) security. However, one must also understand the strength of the assumption required by TEEs: the data owner fully trusts the manufacturer. TEE also requires trusting the protocol implementation because there exist other flaws: side-channel attacks (brasser_software_2017; wang_leaky_2017).
Sasy et al. (sasy_fast_2022) presented an excellent example of the risk of TEE. While some shuffle DP papers used Prochlo (bittau_prochlo_2017) (a TEE-based secure shuffler) as motivation, Sasy et al. (sasy_fast_2022) argued that recent attacks against TEE make Prochlo insecure. Hence, even if Prochlo seemed at first totally secure, it turned out to be insecure in practice due to TEE vulnerabilities.
@@ 8. Conclusion and future directions
Linear complexity for 1-server secure shuffling. Recent works (bittau_prochlo_2017; shi_non-interactive_2021; bunz_non-interactive_2021) approached linear complexity (i.e., $O(n)$ communications and computations) for 1-server secure shuffling, the optimal complexity for 1-server shuffling.
Prochlo has linear complexity but is vulnerable to side-channel attacks. Recent secure TEE-based solutions (ngai_distributed_2024; gu_efficient_2023) avoided these attacks, but have a complexity of $O(n\log n)$. Finally, NIDAR verifies fewer security properties than the rest of the solutions and only has subquadratic costs.
@@
Formalization of the “imperfect” shuffle DP model. Several works (kwon_atom_2017; bunz_non-interactive_2021; langowski_trellis_2023; zhou_power_2022; balle_amplification_2023) proposed relaxed anonymity definitions to conceive more scalable secure shufflers. While some works (zhou_power_2022; balle_amplification_2023) adapted the shuffle DP model to their “imperfect” shuffling, no DP results exist for the relaxation of Langowski et al. (langowski_trellis_2023). Thus, further theoretical work is still needed to formalize these relaxations, and present DP results for shufflers with a relaxed anonymity definition.
Computational shuffle DP. Subsection 7.4 explained why only computational Shuffle DP is practical. While Mironov et al. (mironov_computational_2009) showed that computational DP can reach better accuracy guarantees than information-theoretic DP, no work has yet studied this relaxation in Shuffle DP. Hence, analogous investigations for Shuffle DP would be necessary to take advantage of our impossibility result and build more accurate computations.
New types of shuffling-based private computation. While most existing applications are using secure shufflers to perform a form of data aggregation, recent works (gascon_computationally_2024; wang_beyond_2025) recently opened the way to other privacy-sensitive applications such as information retrieval or taxi hailing could be enabled by secure shufflers. Thus, further research into new shuffling-based private computations is a promising direction. The concept of “private individual computation” introduced by wang_beyond_2025 represents a valuable framework to structure such research effort.